A storage battery contains a solution of H2SO4 38% by weight. At this
concentration van’t Hoff factor is 2.50. At what temperature will the battery
contents freeze?
(kf for water =1.86k kg/ mol)

Grade:upto college level

2 Answers

Archana s
6 years ago
∆T = I *  kf  * m 
∆T = 29.08 
∆T = 273-29.08 =  243.92k
Solubility decrease with temperature
Or
∆T = kf * m
9.3 = 1.86 * 50 / w * 1000 / 62
W = 161.29
When pressure more than osmotic pressure is applied desalination of water 
 
Naman jain
6 years ago
W2=38gW1=62gM2=98gI=2.50T=?Kf for water=1.86∆T=Kfxm=100xkfxW2÷W1xM2=1000×1.86×38÷62×98=11.633i=observed ∆Tf÷calculated∆TfObserved∆Tf=ixcalculated∆Tf=2.50×11.63=29.08Freezing point=273-20.08=243.92